standing wave on a stretched string with a tension force F_T and of length L = 2 m has the following equation: y(x,t) = 0.1 sin⁡(2πx) cos⁡(100πt). How many loops would appear on the string if the velocity is increased by a factor of 2 while the frequency is held constant?


4 loops


12 loops


8 loops


2 loops


6 loops

Clear selection



1
Expert's answer
2021-05-27T14:00:58-0400

Comparing the given equation with standard equation

"y=Asin(kx)cos(\omega t)"

"A=0.1\\k=2\pi\\\omega=100\pi"


"k=2\pi\lambda"

"\lambda=1\space m"

Let velocity = "v"

frequency = "f"


"v=f\times\lambda"

"f=\dfrac{v}{\lambda}"


Let "\lambda'" be the wavelength when velocity is increased by a factor of 2 while frequency is constant

"f=\dfrac{2v}{\lambda'}"


Equating frequencies,

"\dfrac{v}{\lambda}=\dfrac{2v}{\lambda'}"

"\lambda'=2\lambda"

"\lambda'=2\space m"


Number of loops, n

"n=2Lλ'=2\times2\times2=8"